OpenStudy (anonymous):

A car dealer offers you two deals on a car that costs $16,000. Please calculate the monthly payment, given these two payment options the car dealer is offering. Payment Option 1: You can finance the car for 60 months with no interest if you make a $3,000 down payment. Payment Option 2: You can finance the car for 72 months (6 years) with 1% simple annual interest and no down payment. (hint: To calculate simple annual interest, use the formula Interest = Principal * Rate * Time (in years). Add the amount of interest to the price of the car.) Which monthly payment amount is lower? Please exp

OpenStudy (anonymous):

part1 3000 down payment means 13000 is principle amount now monthly payment would be 13000/6o=??? if there is no interest part 2 now principle=16000 interest=1 72 months means 6 years so interest = principle*1%*6years= (16000/100)*1*6=960 so total amount=16000+960=16960 now monthly payment=16960/72=??? now calculate and compare both results to get answer